OpenStudy (anonymous):

simplify cos 0/ tan 0 cot0

OpenStudy (anonymous):

the options are a)Sin Theta B)Cos Theta c)Csc Theta D)Sec Theta Pleaseeee help me :(

OpenStudy (ankitshaw):

give it a try its simple...

OpenStudy (anonymous):

I did, I have to have a whole semester worth the online schooling done by today or I don't graduate. I need all the help I can get right now.. lol

OpenStudy (akashdeepdeb):

Hey @tayvaught14 So, to solve this question, you need to have an understanding of Trigonometric Identities. This is what the question asks. \[\frac{cos ~\theta}{tan~\theta . cot~\theta}\] Now what we know is: \[tan ~\theta = \frac{1}{cot~\theta}\] So, you can write, tan \(\theta\) as that and then see what you get! :D Try now. :)

OpenStudy (anonymous):

Okay hold on, thank you :)

OpenStudy (anonymous):

Wait so do I put the fraction for tan theta in the other fraction? I'm sorry I'm very confused on all of this.

OpenStudy (akashdeepdeb):

No problem! I'll help. These are some trigonometric Identities: \[\sin \theta = \frac{1}{csc \theta}\]\[cos \theta = \frac{1}{sec\theta}\]\[tan\theta = \frac{1}{\cot\theta}\] Are you familiar with these?

OpenStudy (anonymous):

Sort of yes

OpenStudy (akashdeepdeb):

Do you know what trigonometric identities represent? They are simple ratios of a right angled triangle:  Let's move on to the more lovable part of trigonometry. Simplification of ratios. In your question, this is what you are asked to find: \[\frac{cos\theta}{tan\theta.cot\theta}\] Try to use the IDENTITIES/ THE RELATIONS BETWEEN THE RECIPROCALS. HERE. Ask if you don't get it. I'll help. But just try once. :D

OpenStudy (anonymous):

Okay, Thank you so much! I will ask you when I think I've figured it out :)

OpenStudy (anonymous):

Is there an equation or something that would help me figure out what goes where. I understand what you sent me, in a sense, but is there a certain way I plug it into the original problem?

OpenStudy (akashdeepdeb):

Yes. Totally! :D This is what you are given: \[\frac{cos\theta}{tan\theta.cot\theta}\] What do you know about \(tan\theta\) from the identities? :D We know that: \[tan\theta = \frac{1}{cot\theta}\] Let us substitute this, in place of \(tan\theta\) in the original equation! We get:\[\frac{cos\theta}{\frac{1}{cot\theta}.cot\theta}\] Okay, we have something now! :D Let's see the denominator. We have \[\cot\theta.\frac{1}{cot\theta}\] which is equal to ? Well, it is equal to 1 ! :D It's just like: \[2 * \frac{1}{2} = 1 = a * \frac{1}{a} = (anything) * \frac{1}{(anything)} = 1\] So we now have our denominator = 1. Numerator still is \(cos\theta\) Thus the fraction now is = \[\frac{cos\theta}{1} = cos\theta\] Getting this? If not, Ask, I'll explain. :)

OpenStudy (anonymous):

so you substitute the fraction for tan in, and then you multiply it by the denominator to get rid of the fraction?

OpenStudy (akashdeepdeb):

Yes, I multiply cot [which was already in the denominator before] and then when I multiply it, I get 1. :)

OpenStudy (anonymous):

I think I understand that! Would you help me with another now that I sort of get the basics?

OpenStudy (akashdeepdeb):

Sure. Try it yourself once. Use the above identities. Go through this link: http://www.mathsisfun.com/algebra/trigonometric-identities.html I'll be around.

OpenStudy (anonymous):

So that page is very helpful thank you! But would you be able to help me get started on this question and I'll finish it? Thank you so much for helping @AkashdeepDeb

OpenStudy (akashdeepdeb):

I already posted the answer. Check it again! ^^

OpenStudy (anonymous):

You don't understand how much you're helping me right now! lol thank you @AkashdeepDeb

OpenStudy (akashdeepdeb):

Glad. :)